A tank with a horizontal sectional area constant at 10 square meter and 4 m high contains water to a depth of 3.5 m. the tank has a circular orifice 5 cm in diameter and located at its side 0.5 m above the bottom. if the coefficient of discharge of the orifice is 0.60, find the duration of flow though the orifice.

1
Expert's answer
2022-01-25T04:43:52-0500

Given "A=10 m^2, h_1=3.5m, h_2=3.5m-0.5m=3m,"

"r=5cm\/2=0.025cm,C_d=0.60"

After time "dt"


"Qdt=-Adh"

"C_d(\pi r^2)\sqrt{2gh}dt=-Adh"

"dt=-\dfrac{A}{\pi C_dr^2\sqrt{2g}}h^{-1/2}dh"

Integrate


"\displaystyle\int_{0}^{T}dt=-\displaystyle\int_{h_1}^{h_2}\dfrac{A}{\pi C_dr^2\sqrt{2g}}h^{-1/2}dh"

"T=-\dfrac{A}{\pi C_dr^2\sqrt{2g}}[2\sqrt{h}]\begin{matrix}
 h_2 \\
 h_1
\end{matrix}"

"T=\dfrac{2A(\sqrt{h_1}-\sqrt{h_2})}{\pi C_dr^2\sqrt{2g}}"

"T=\dfrac{2(10m^2)(\sqrt{3.5m}-\sqrt{3m})}{\pi (0.6)(0.025m)^2\sqrt{2(9.81m/s^2)}}"

"=532 s"

The duration of flow though the orifice is "532s."